OpenStudy (anonymous):

How do i find the two numbers that x equals from 3sin(x)=1?

OpenStudy (anonymous):

Well, you want to isolate the sin(x) and get sin(x) = 1/3. Then, what angle measures \[0\le x<2\pi\] will make sin(x) equal 1/3?

OpenStudy (anonymous):

I'm not sure how to figure that out either

OpenStudy (anonymous):

Well, to my knowledge, there aren't closed real solutions to the equation. I think you would have to just do arcsin(1/3) on a calculator. What class are you taking?

OpenStudy (anonymous):

ok ill try that out and im taking Calculus

OpenStudy (anonymous):

Oh. Are you doing newton's method, then?

OpenStudy (anonymous):

no its part of a question to find the area between two curves

OpenStudy (anonymous):

Oh, okay. And you're trying to find the intersection of y = 3sin(x) and y = 1? You may just want to compute the intersection on a graphing calculator or something.